OpenStudy (anonymous):

Which of the following points is in the solution set of y < x2 - 2x - 8? (-2, -1) (0, -2) (4, 0)

OpenStudy (anonymous):

please answer i need help!!!!

OpenStudy (anonymous):

-1<4+4-8=0 -1<0 true, so first point is good. repeat the same for other 2 points

OpenStudy (anonymous):

-2<0^2-0-8=-8 -2<-8 is not a true statement , so second point is not a solution

OpenStudy (anonymous):

0<4^2-2*4-8=16-8-8=0 0<0 is not a true statement , should be 0=0 so your only solution is point: (-2, -1)